OpenStudy (anonymous):

$2.99 for socks with 5.5% as the sales tax

jigglypuff314 (jigglypuff314):

5.5% = 0.055 total amount = original price + tax so 2.99 + (0.055)(2.99) = ?

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\%\ of \ b=\cfrac{{\color{red}{ a}}}{100}\times b\qquad \qquad {\color{red}{ 5.5}}\%\ of \ 2.99=\cfrac{{\color{red}{ 5.5}}}{100}\times 2.99\)

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

still not getting it

jigglypuff314 (jigglypuff314):

plug in 0.055 times 2.99 plus 2.99 into your calculator

OpenStudy (anonymous):

i got 3.15445

jigglypuff314 (jigglypuff314):

yep :) then round to the hundredths place :) 3.15445 ^ hundredths place

OpenStudy (anonymous):

oh thank you so much

jigglypuff314 (jigglypuff314):

glad we could help :)